OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–2, 15), (9, –18)

OpenStudy (anonymous):

Let p(x,y) be a point on the line. Equate the slopes calculated using p and the given points. \[\frac{y-15}{x+2}=\frac{y+18}{x-9}\]Solve the above for y.\[y=-3x+9 \]